# ActiView: Evaluating Active Perception Ability for Multimodal Large Language Models

### v1

@@ 5 Results and Analysis
Analysis of Cases
We demonstrate three cases for the three proposed pipeline in Figure 5, and provide additional examples of integrating human-annotated visual clues hints in Appendix G. These results are generated by GPT-4 models and Gemini-1.5 models. Case (a) stands for the zooming evaluation, where models successfully identify the view containing useful information and generate the correct result. Case (b) illustrates a failure in the shifting-R evaluation, where all the models continue shifting to new views until all views are used. Though including the correct views, the additional views severely distract the reasoning process, where three out of four employed models produce incorrect answers. To explore how human-like mixed evaluation affects the visual reasoning process, we further exam this failure case using GPT-4o. As shown in Figure 5 case (c), GPT-4o first zooms into the “upper left” and “upper right” views, then discards the “upper right” view and shifts to the “lower left” one, which finally leads to the correct answer. Notably, in the final preserved views, distracting information (the highest price tag on papaya, “69”) is screened out. This indicates that GPT-4o exhibits decent active perception abilities to move the field of view, locate details, and filter out distracting information
6 Conclusion and Future Work
@@
In this study, we assess the active perception abilities of models through question answering. While this approach presents significant challenges for current MLLMs, it does not encompass all aspects of active perception. For instance, techniques like tool learning and multi-agent collaboration could also be investigated concerning the factors such as perspective distortion, multi-sensor integration, and the incorporation of more dynamic or interactive environments. Given that these limitations exceed the scope of a single paper, we will leave them as future work.
Appendix A Data Collection and Annotation
@@ Appendix B Human Evaluation
Appendix C Models
Our evaluated APIs include GPT-4V (OpenAI, 2023), GPT-4o (OpenAI, 2024), and Gemini-1.5-pro (Reid et al., 2024). For open-source models, we include GLM-4V (Du et al., 2022), SEAL (Wu & Xie, 2023), Mini-Gemini (Li et al., 2024c), MMICL (Zhao et al., 2024), InternVL series (Chen et al., 2023; 2024), Brote (Wang et al., 2024b), LLaVA-1.6 (Liu et al., 2024), LLaVA-OneVision (Li et al., 2024a), mPLUG-Owl series (Ye et al., 2024b; a), MiniCPM-Llama3-V-2.5, MiniCPM-V 2.6, Qwen2-VL (Wang et al., 2024a). Details of these models are listed in Table 6. Considering models of different scales, we include a total of 27 models.
@@ Appendix C Models
InternVL-Vicuna-7B (Chen et al., 2023) Vicuna-7B InternViT
InternVL-Vicuna-13B (Chen et al., 2023) Vicuna-13B InternViT
InternVL-Vicuna-13B-448px (Chen et al., 2023) Vicuna-13B InternViT‑300M‑448px
InternVL2-8B (Chen et al., 2024) internlm2_5‑7b‑chat InternViT‑300M‑448px
@@
MMICL-XXL-11B (Zhao et al., 2024) FlanT5-XXL-11B EVA-G
Brote-IM-XXL-11B (Wang et al., 2024b) FlanT5-XXL-11B EVA-G
MMICL-XL-3B (Zhao et al., 2024) FlanT5-XL-3B EVA-G
Brote-IM-XL-3B (Wang et al., 2024b) FlanT5-XXL-3B EVA-G
@@
Appendix D Calculation of View Selection Recall
@@
We follow the recall metric to evaluate the performance of the view selection for zooming setting and the missing view examination for shifting settings. We denote the selected views containing human-annotated clues as $TP_{op}$, where $op$ refers to either “zoom”, “shift” or “mix”. $FN_{op}$ refers to views that contain human-annotated clues but are not selected for answering questions. Finally, the $R_{\text{select}}$ is calculated as follows:
@@
The view selection recalls of selected models are listed in Table 4.
Table 7: The evaluation of active perception abilities on our benchmark, including zooming (for limited resolution scenarios), and shifting (for scenarios of limiting the field of views). “Model AVG”: average scores of column “Zooming”, “Shifting-R”, “Shifting-E”, “Shifting-M”, and “Shifting-H”. The best scores of each column are bolded and the best scores in each model types are highlighted.
@@
Full image Zooming Single View Shifting-R Shifting-E Shifting-M Shifting-H
@@
Table 8: Results on sub-classes of zooming evaluation.
@@
Table 9: Results on each sub-classes of Shifting-R, shifting with random initial views.
@@
Table 10: Results on sub-classes of Shifting-E (the easy-level shifting evaluation), where initial views contain clues for answering the question.
@@
Table 11: Results on each sub-classes of Shifting-M (the medium-level shifting evaluation), where initial views contain only partial clues for answering the questions.
@@
Table 12: Results on each sub-classes of Shifting-H (the hard-level shifting evaluation), where initial views do not display clues for answering the questions. Models should decide whether to shift to the next view all by themselves.
@@
Appendix E Experimental Results
We reported the full results of 27 models in Table 7. This table preserves the conclusions as discussed in by Table 2. The detailed results of each categories are listed in Table 8, Table 9, Table 10, Table 11, and Table 12, for zooming, shifting-R, shifting-E, shifting-M, and shifting-H, respectively.
E.1 Analysis of Results on Zooming Evaluation
@@
E.2 Analysis of Results on Shifting Evaluation
@@
Results of shifting-R evaluation are shown in Table 9, and the level-specified shifting evaluation are listed in Table 10, Table 11 and Table 12. Similar to that of zooming evaluation, results on environment-centric category are significantly better than the ones on object-centric and event-centric categories. The results of proprietary models are better than the results of open-source models, and that models for multiple images perform better than models for single image. We observe a trend where, as the difficulty increases, the superiority of open-source multi-image models becomes more evident.
@@
There is overall trend for all the sub-classed that the accuracy decreases as the difficulty is getting increased. However, exceptional performances are identified for Gemini-1.5-pro, Idefics3-8B-Llama3, and Mini-Gemini-7B-HD, where the results of shifting-H even outperform the results of shifting-E. There is a potential reason that the hard-level evaluation starts with less relevant views and appends more useful views at the end of the image sequence. The performance of these models are more significantly influenced by the order of presented images compared to the rest models. The degradation of performance is more remarkable for the environment-centric and the object-centric categories compared to the event-centric category. Regarding the increasing of difficulty for the environment-centric and the object-centric categories, we observe gaps of about 10% for models such as LLaVA-OneVision, Idefics2-8B, Brote, MMICL, GLM-4V-9B and Mini-Gemini-7B. These observations indicate that different initial perceptual fields have distinct impacts on instances that requiring demanding attention on subtle changes of fine-grained objects. Results show that GPT-4o consistently outperforms other models in the average score of the environment-centric category, implying robust event capture and understanding capabilities in multi-image scenarios.
Appendix F Discussion on Image Splitting and Processing Strategies
F.1 Image Splitting Settings
Table 13: Experimental results of different splits.
Model Splits Zooming Shifting-R AVG
LLaVA-1.6 7B 4 68.92 51.69 60.31
LLaVA-1.6 7B 6 73.23 53.85 63.54
LLaVA-1.6 7B 8 72.92 48.61 60.77
LLaVA-1.6 7B 9 66.46 46.16 56.31
LLaVA-1.6 7B 16 69.23 46.15 57.69
LLaVA-1.6 13B 4 65.23 53.85 59.54
LLaVA-1.6 13B 6 71.69 46.46 59.07
LLaVA-1.6 13B 8 71.84 44.00 57.92
LLaVA-1.6 13B 9 72.00 43.69 57.84
LLaVA-1.6 13B 16 73.31 43.23 58.27
F.2 Strategies of Processing Multiple Images for Single-image Models
Model Visual Info. Type Zooming Shifting-R
LLaVA-1.6 7B Image concatenation 68.92 51.69
Textual descriptions 60.31 -8.61 53.83 +2.14
LLaVA-1.6 13B Image concatenation 65.23 45.85
Textual descriptions 60.00 -5.23 43.69 -2.16
mPLUG-Owl2 7B Image concatenation 55.38 47.38
Textual descriptions 62.77 +7.39 54.15 +6.77
MiniCPM-Llama3-V-2.5 Image concatenation 61.25 60.92
Textual descriptions 61.25 -0 60.31 -0.61
Appendix G Case Study of Providing Human-annotated Clues

### v2

@@ 5.3 Analysis of View Processing Strategies
6 Conclusion
@@
Limitations
In this study, we utilize the form of VQA and intermediate reasoning behaviors to assess active perception abilities of models. While pipeline presents significant challenges for current multimodal language models (MLLMs), it does not encompass all aspects of active perception. For instance, it overlooks factors such as perspective distortion, multi-sensor integration, the incorporation of more dynamic or interactive environments, and real-time manipulation. With the development of reasoning ability of models such as OpenAI o1, o3 and DeepSeek-R1, we believe the active perception ability will also emerge or be improved. Moreover, techniques like tool learning and multi-agent collaboration could potentially enhance active perception performance based on existing MLLMs, making these areas worthy for future exploration and improvement. Additionally, active perception could also be used to improve other challenging topics such as hallucination and real-life application such as auto-driving. Also, Considering that these exceed the scope of a single conference paper, we do not include them in this paper, and we solely evaluates the inherent active perception capabilities of the MLLMs, aiming at an in-depth investigation of our primary focus.
Appendix A Data Details
@@ A.2 Rules for Annotation
  • Clues: regions in the image that contribute to your annotated answer.
A.3 Detailed Category Description
When perceiving an image, humans intuitively focus on three principle aspects: the environment depicted in the image, the primary objects, and the event that these objects are engaged in. Correspondingly, we summarize the questions in our benchmark into three main categories, environment-centric (Type I), object-centric (Type II), and event-centric (Type III) categories, which are further divided into eight sub-classes according to the specific type of visual information and visual features used for answering the questions, as displayed in Figure 2. For the environment-centric category, three sub-classes are developed:
  • Geo-Localization (Geo-Loc) focuses on geographical features that are unique to a country or a city, and requires models to identify geographical locations depicted in target images. Typical questions are “Where is this place located?”,“In which country is the photo taken?”, and etc. Images in this class usually contain unique landmarks such as the Eiffel Tower in Paris and the Atomium in Brussels.
  • Orientation (Orient) challenges models to exploit natural orientation information for answering the questions, such as the position of shadows, the position of the sun, and the directional information on street signs. Questions of this type include “Is this a sunset or a sunrise?”, “Where is the sunlight coming from?” and etc.
  • Daily-location (Daily-Loc). To distinguish from Geo-Loc, this sub-class concentrates on locations in everyday life that could appear in most of the cities and are not unique to a certain city or country. Images in this sub-class usually depict scenes of museums, restaurants, shops, etc. The corresponding questions include “Where is this picture most likely taken?”, “Is there a music school nearby?”, and etc.
  • Object-attribute (Obj-Attr) addresses objects attributes while distracting information, that potentially lead to incorrect answers, appears in the images. Shown by the Obj-Attr case in Figure 2, the highest price, 69 per kilogram, corresponds to papaya rather than watermelon.
  • Object-relation (Obj-Rel) concentrates on the spatial relationships among multiple objects, while the questions do not directly ask for the spatial relationships. It requires models to reason for the correct answer via spatial information. Figure 2 displays an Obj-Rel case in which models should be aware of the relative positions of the feet of the person to the water.
  • Counting (Count-Dis). Although it focuses on the number of objects, different from the counting tasks in other benchmarks (Fu et al., 2023; Yu et al., 2023), there are similar but distracting information about the targets in our images. These distracting objects easily confuse models and challenge the abilities to understand and strictly follow instructions. As the Count-Dis case in Figure 2, the jerky on the table are distracting to the answer of question “How many pieces of jerky are hanging on the wall?”.
  • Event-single (Event-S). There is only one item or person involved in the target event. For example, the image for Event-S in Figure 2 shows one person driving without other people presenting in the image.
  • Event-multi (Event-M). Different from Event-S, events of this type happen among multiple items or people. In the Event-M case in Figure 2, the “woman in blue” is engaged in a photo shooting activity in which she is posing and another person is taking photo for her. It requires models to distinguish the event or events that each entities are engaged in.
@@ Appendix B Human Evaluation
Model Claude GPT-4o Qwen2-VL MiniCPM-V 2.6 Idefics3 Brote-IM-XL
ACC 2.14 2.45 23.38 26.77 44.92 40.00
ACC(guess) 26.07 37.73 42.77 41.54 47.38 40.00
Table 7: Results of text-only evaluation. ACC: answer with commonsense only without random guessing. ACC(guess): guess the answer according to commonsense.
@@
B.1 Analysis of Human Performances
We assess the difficulty and reliability of our benchmark upon human performances in Table 6. We employ two settings for human evaluation, where “Human” asks annotators to answer questions only by themselves, and “Human*” allows annotators to use the Internet and LLMs for extra knowledge that could help answer the questions. This “Human*” evaluation aims at fair comparison as humans may not be exposed to knowledge that never appear in their everyday life, while most of MLLMs should be aware of these knowledge from the training data. Human presents an average accuracy of 77.53%, suggesting that our benchmark is challenging even for human. For a fair comparison with large models, “Human*” achieves an average of 84.67% by allowing searching for world knowledge from the Internet or using LLMs. Human performances (84.67%) are more than doubled of the random result (33.95%), while some models present only slightly higher accuracies compared to the random result. These indicate the potential for models to get improved.
@@ B.1 Analysis of Human Performances
SEAL (Wu and Xie, 2023) Vicuna-7B CLIP ViT-L/14
InternVL-Vicuna-7B (Chen et al., 2023) Vicuna-7B InternViT
InternVL-Vicuna-13B (Chen et al., 2023) Vicuna-13B InternViT
InternVL-Vicuna-13B-448px (Chen et al., 2023) Vicuna-13B InternViT-300M-448px
InternVL2-8B (Chen et al., 2024) internlm2_5-7b-chat InternViT-300M-448px
@@
Qwen2.5-VL-3B (Team, 2025) Qwen2.5 trained from scratch
Qwen2.5-VL-7B (Team, 2025) Qwen2.5 trained from scratch
Deepseek-VL-7B (Lu et al., 2024) Deepseek Siglip-large-patch16-384
Deepseek-VL2 (Wu et al., 2024) Deepseek2 Siglip-400m
@@
MMICL-XXL (Zhao et al., 2024) FlanT5-XXL-11B EVA-G
Brote-IM-XXL (Wang et al., 2024b) FlanT5-XXL-11B EVA-G
MMICL-XL (Zhao et al., 2024) FlanT5-XL-3B EVA-G
Brote-IM-XL (Wang et al., 2024b) FlanT5-XL-3B EVA-G
@@
Appendix C Text-only Evaluation
We provide a text-only evaluation to measure the amount of commonsense answers with providing images in our benchmark. We conducted two experiments:
  • Commonsense-only evaluation. This evaluation aims at measuring the amount of questions that can be answered only via commonsense knowledge without searching for visual clues in the image. The template is as follows: “Please answer questions based on you commonsense knowledge. If you are not able to answer the question based soly on the commonsense knowledge you’ve acquired, please response with ‘None’. Question Options Your answer:”
  • Commonsense and data bias evaluation. Considering that current models are trained with a large amount of data and various tasks, they could potentially memories the most frequent answers given a image-question pair. We implement another template to evaluate the amount of data that can be correctly guessed without corresponding context. The template is as follows: “Please answer questions based on you commonsense knowledge. If you are not able to answer, please select a most probable one. Question Options Your answer:”
Results for these text-only evaluations are listed in Table 7. This table indicates that questions in our benchmark cannot be simply answered via commonsense, where two powerful models GPT-4o and Claude achieves only 2.45% and 2.14% for commonsense-only evaluation. The row of ACC(guess) presents results of generating the most probable answers, reflecting the bias obtained from the training corpus. The differences between these two type of evaluation are caused by the ability of instruction-following. We found that Idefics3 and Brote-IM-XL present weaker instruction-following ability compared to other models in this table, that they still exhibit a behavior of guessing when commonsense cannot be used to answer the questions.
Overall, our benchmark requires elaborate observation of the given images and comprehensive understanding of image-question pairs, which cannot be solved simply by commonsense.
Appendix D Models
We investigate both proprietary and open-source models. The proprietary models include widely discussed GPT-4o (OpenAI, 2024), Gemini-1.5-pro (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024). For open-source models, we carefully select recent and commonly used models of different structures and of difference scales, such as model families of MiniCPM-V Yao et al. (2024), LLaVA (Liu et al., 2023b, c), mPLUG-Owl (Ye et al., 2024b, a), Idefics (Laurençon et al., 2024, 2024), and etc. Since the awareness of fine-grained details and instruction-aware visual features are significant indicators during evaluation, we also include models specifically optimised on these aspects, such as SEAL (Wu and Xie, 2023) for fine-grained details understanding, and Brote (Wang et al., 2024b) which is trained from InstructBLIP (Dai et al., 2023) for instruction-aware and multi-image comprehension. Details of these models are listed in Table 8. Considering models of different scales, we include a total of 27 models. These models are divided into two types, single-image models that accepting only one image per input, such as LLaVA-1.6 (Liu et al., 2023b) and MiniCPM-Llama3-V-2.5 (Yao et al., 2024); and multi-image models that allow more than one images to appear in the same input, such as Brote and Idefics. We describe the approaches for integrating multiple views into the input for the two types of models in Appendix F.2 and Appendix F.3.
Appendix E Evaluation Pipelines
This section will discuss motivations and settings of each pipelines in detail.
Zooming pipeline.
It focuses on one of the fundamental factors, zooming, and evaluates the ability to locate and determine fine-grained information necessary to answer questions. As illustrated in Figure 4 (a), this pipeline contains two stages, the view selection and the question answering stages. To simulate the zooming operation, models are required to first select sub-views to be zoomed given the initial view, then answer questions based on these zoomed views. The initial view used in this pipeline is the full image with size $w\times h$. Each of the selected sub-views will be resized to size $w\times h$, the same as the initial view. In Figure 4 (a), the zoomed right-upper view is resized as a $w\times h$ image, and so does the zoomed left-lower view. Afterwards, models answer the question given the two zoomed views. Please refer to Appendix J.2 for prompt templates.
Shifting pipeline.
It addresses the other fundamental factor, shifting, and emphasizes the ability to navigate perceptual fields incrementally, mimicking real-world scenarios where full context is unavailable. It evaluates the ability to shift perceptual fields for missing information and to deduce the answer given perceived perceptual fields following templates in Appendix J.3. This is also a two-stage pipeline as in Figure 4 (b). To simulate the movement of human eyes, models are presented with an initial view, size $w\times h$, which is a cropped field from the original image, and are asked to determine if the current views are sufficient for answering. Upon receiving positive responses, models are prompted to produce answer given the current view or views. If the model requires more views to infer the answer, an adjacent view will be given until the model can answer the question. For this pipeline, we further assign different difficulties according to human-annotated visual clues contained in the initial views as follows:
  • Shift-R: randomly selected initial views.
  • Shift-E: easy-level evaluation, where initial views contain at least one entire visual clue for answering the question.
  • Shift-M: medium-level evaluation, where initial views contain only partial visual clues for answering the question.
  • Shift-H: hard-level evaluation, where no visual clues appear in the initial views.
Mixed pipeline.
While the above pipelines permit either zooming or shifting individually, we also implement an automated mixed setting that does not specify the type of active perception ability required. As illustrated in Figure 4 (c), models must independently decide whether to zoom and/or shift to different perceptual fields. Unlike the zooming pipeline, where the model answers questions based on all selected views, in the mixed pipeline, a view would be discarded after selection if the model recognizes it as irrelevant to the question. Compared to the shifting pipeline, the mixed pipeline also provides access to the full image view in addition to cropped sub-views. Appendix J.4 records the employed prompt templates. This pipeline requires models to account for all the sub-views and the full image for unbiased operation determination and view selection. Otherwise, it is at risk of reverting to zooming or shifting evaluation without sufficient and unconverted visual information. Therefore, the mixed pipeline emphasizes the autonomy of models and is only applied to multi-image models.
Appendix F Discussion on Image Splitting and Processing Strategies
F.1 Image Splitting Settings
Model Splits Zooming Shift-R AVG
LLaVA-1.6 7B 4 68.92 51.69 60.31
LLaVA-1.6 7B 6 73.23 53.85 63.54
LLaVA-1.6 7B 8 72.92 48.61 60.77
LLaVA-1.6 7B 9 66.46 46.16 56.31
LLaVA-1.6 7B 16 69.23 46.15 57.69
LLaVA-1.6 13B 4 65.23 53.85 59.54
LLaVA-1.6 13B 6 71.69 46.46 59.07
LLaVA-1.6 13B 8 71.84 44.00 57.92
LLaVA-1.6 13B 9 72.00 43.69 57.84
LLaVA-1.6 13B 16 73.31 43.23 58.27
Table 9: Experimental results of different splits.
Model Visual Info. Type Zooming Shift-R
LLaVA-1.6 7B Image concatenation 68.92 51.69
Textual descriptions 60.31 -8.61 53.83 +2.14
LLaVA-1.6 13B Image concatenation 65.23 45.85
Textual descriptions 60.00 -5.23 43.69 -2.16
mPLUG-Owl2 7B Image concatenation 55.38 47.38
Textual descriptions 62.77 +7.39 54.15 +6.77
MiniCPM-Llama3-V-2.5 Image concatenation 61.25 60.92
Textual descriptions 61.25 -0 60.31 -0.61
F.2 Processing of Views
The question answering stage of the zooming, shifting, and mixed pipelines, as well as the missing view examination stage of the shifting pipeline, require multi-image inputs if multiple views are selected. In this paper, we primarily focus on the interleaved multi-image setting, since it is more practical and natural compared to the single-image setting. Multi-image models can naturally read and understand multiple views at one time (in the form of different images) during evaluating, and we directly format the images and text in an interleaved format. However, we also propose methods for evaluating powerful single-image models. For these models, we employ two strategies to enable simultaneous understanding of different views. One is to concatenate the required views into a single flattened image, and the other preserves merely the current view as an image while converting the remainings into textual descriptions. The following subsection discusses this in detail.
F.3 Strategies of Processing Multiple Images for Single-image Models
Appendix G Measurements of View Selection
@@
We follow the recall, precision and F1 metrics to evaluate the performance of the view selection for zooming setting and the missing view examination for shifting settings. We denote the selected views containing human-annotated clues as $TP_{op}$, where $op$ refers to either “zoom”, “shift” or “mix”. $FN_{op}$ refers to views that contain human-annotated clues but are not selected for answering questions, and $FP_{op}$ refers to views selected but do not contain human-annotated clues. Finally, the precision, $P_{\text{select}}$, is calculated as follows:
$P_{\text{select}}=\frac{TP_{op}}{TP_{op}+FP_{op}},{\quad op\in\{\text{zoom},\text{shift},\text{mix}\}}.$ (1)
It measures the proportion of selected views that are actually relevant. It reflects the model ability to avoid unnecessary or irrelevant views. A higher precision indicates that the model is more efficient in identifying only the information necessary for answering the question. The recall, $R_{\text{select}}$, is calculated as follows:
@@
This recall score measures the proportion of views correctly identified by the model out of all views containing human-annotated clues. It reflects the model ability to capture the required information. A higher recall indicates that the model is less likely to miss important views during the zooming process. Accordingly, F1 score of view selection, $F_{1}$ is computed as:
$F_{1}=\frac{2\cdot P_{\text{select}}\cdot R_{\text{select}}}{P_{\text{select}}+R_{\text{select}}}$ (3)
@@
Full image Zooming Single View Shift-R Shift-E Shift-M Shift-H
@@
Table 11: The evaluation of active perception abilities on our benchmark, including zooming (for limited resolution scenarios), and shifting (for scenarios of limiting the field of views). “Model AVG”: average scores of column “Zooming”, “Shift-R”, “Shift-E”, “Shift-M”, and “Shift-H”. The best scores of each column are bolded and the best scores in each model types are highlighted.
Appendix H Experimental Results
We reported the full results of 27 models in Table 11. This table preserves the conclusions as discussed in by Table 2. The detailed results of each categories are listed in Table 12, Table 13, Table 14, Table 15, and Table 16, for zooming, Shift-R, Shift-E, Shift-M, and Shift-H, respectively.
@@ Appendix H Experimental Results
Table 12: Results on all sub-classes of zooming evaluation.
@@
Table 13: Results on each sub-classes of Shift-R, shifting with random initial views.
@@
Table 14: Results on sub-classes of Shift-E (the easy-level shifting evaluation), where initial views contain clues for answering the question.
@@
Table 15: Results on each sub-classes of Shift-M (the medium-level shifting evaluation), where initial views contain only partial clues for answering the questions.
@@
Table 16: Results on each sub-classes of Shift-H (the hard-level shifting evaluation), where initial views do not display clues for answering the questions. Models should decide whether to shift to the next view all by themselves.
H.1 Analysis of Results on Zooming Evaluation
@@
H.2 Analysis of Results on Shifting Evaluation
@@
The shifting pipeline aims at mimicking the scenario when humans look for more visual information by shifting the perceptual fields, the previously perceived views cannot be simply erased from memory, and new views are integrated incrementally. Results of Shift-R evaluation are shown in Table 13, and the level-specified shifting evaluation are listed in Table 14, Table 15 and Table 16. Similar to that of zooming evaluation, results on environment-centric category are significantly better than the ones on object-centric and event-centric categories. The results of proprietary models are better than the results of open-source models, and that models for multiple images perform better than models for single image. We observe a trend where, as the difficulty increases, the superiority of open-source multi-image models becomes more evident.
@@
There is an overall trend for all the sub-classed that the accuracy decreases as the difficulty is getting increased. As shown by typical results for LLaVA-OneVision and GLM-4V-9B in Table 2, the gaps between Shift-E and Shift-H are as large as 8.61% and 7.70%, respectively. However, exceptional performances are identified for Gemini-1.5-pro, Idefics3-8B-Llama3, and Mini-Gemini-7B-HD, where the results of Shift-H even outperform the results of Shift-E. One of the reason could be the recall of selected views. For Gemini-1.5-pro, the recall for Shift-H is 47.73%, over 1 point higher than Shift-E (45.29%). We conclude that Gemini-1.5-pro achieves higher accuracy on Shift-H due to the acquisition of more proper views. While Idefics3 presents a different trend. It maintains a recall of 74.64% from Shift-E to Shift-H, but achieves a higher accuracy on Shift-H. There is another potential reason that the performance gain of this model comes from the order of input views. The hard-level evaluation starts with less relevant views and appends more useful views at the end of the image sequence, and the performance of these models are more significantly influenced by the order of presented images compared to the rest models. The degradation of performance is more remarkable for the environment-centric and the object-centric categories compared to the event-centric category. Regarding the increasing of difficulty for the environment-centric and the object-centric categories, we observe gaps of about 10% for models such as LLaVA-OneVision, Idefics2-8B, Brote, MMICL, GLM-4V-9B and Mini-Gemini-7B. These observations indicate that different initial perceptual fields have distinct impacts on instances that requiring demanding attention on subtle changes of fine-grained objects. Results show that GPT- 4o consistently outperforms other models in the average score of the environment-centric category, implying robust event capture and understanding capabilities in multi-image scenarios.
H.3 Analysis of View Selection
Our evaluation pipelines involve selecting useful view in their first stages. The reliability of selected views plays a crucial role in the following question answering stage. We compute the recall of used views following Equation 2 in Appendix G, and include results in Table 17, along with accuracy of providing models with groundtruth views. Overall, lower selection recall tends to correlate with lower question answering accuracy. For example, Idefics3-8B-Llama3 and InternVL2-8B present the lowest recalls (41.09%) among multi-image models in Table 17, leading to lower zooming evaluation accuracies of 56.00% and 58.15%, respectively.
For zooming evaluation, we also investigate the performance when the given groundtruth views that contain human-annotated clues. Generally, models are prompted to generate more accurate answers compared to the pure zooming setting. However, mPLUG-Owl3, Gemini-1.5-pro, and LLaVA-OneVision are only exceptional, whose performance slightly degrade when given the visual clues. We argue that these models are better at the question answering task rather than exhibiting active perception ability. Additionally, we observe that shifting evaluations tend to require more views to be used for answering questions than zooming evaluations, yet it often results in inferior overall performance compared to zooming. For the shifting evaluation, some models keep shifting view until all four views are inquired. However, this does not necessarily support a better accuracy, as some views contain redundant information that might distract the model during reasoning. Additionally, we observe shifting evaluations tend to require more views to be used for answering questions than zooming evaluations, yet it often results in inferior overall performance compared to zooming. This is because some of the current advanced models struggle to either move their field of views for necessary visual details, or screen out distracting information. Therefore, we believe that more attention should be paid to evaluating and enhancing active perception abilities of MLLMs given constraint perceptual fields.
[…]
Multi-image Models
GPT-4o 73.54 68.62 69.03 2.29 67.08 60.54 3.26
mPLUG-Owl3 60.62 60.92 68.57 2.66 51.69 74.62 4.00
Claude 3.5 Sonnet 72.31 71.69 67.64 2.19 65.23 45.52 2.47
Qwen2-VL 65.85 64.62 64.61 2.35 61.23 74.62 4.00
Gemini-1.5-pro 72.00 72.31 62.63 2.10 67.08 46.33 2.46
MiniCPM-V 2.6 62.77 61.85 57.03 2.20 54.77 54.83 2.98
LLaVA-OneVision 64.92 65.23 46.67 2.35 53.54 37.14 2.02
Idefics3-8B-Llama3 60.92 58.15 41.09 1.52 61.85 74.62 4.00
InternVL2-8B 73.23 56.00 41.09 1.53 54.77 45.75 2.61
Single-image Models
InternVL-Vicuna-13B 68.00 62.77 83.47 3.31 53.85 69.73 3.75
LLaVA-1.6 13B 67.69 68.92 68.57 2.65 51.69 74.62 4.00
SEAL 56.92 54.77 68.22 2.74 42.15 71.48 3.83
MiniCPM-Llama3-V-2.5 62.20 61.25 66.12 2.46 53.85 63.56 3.42
mPLUG-Owl2-7B 67.38 55.38 47.61 1.97 47.38 74.62 4.00
GLM-4V-9B 74.46 56.92 30.62 1.08 56.92 50.17 2.64
[…]
Appendix I Case Study
In this section, we demonstrate three cases for the three proposed pipelines in Figure 5, and provide additional examples of integrating human-annotated visual clues hints in Figure 6.
I.1 Analysis of Cases from Each Pipelines
These results are generated by GPT-4 models and Gemini-1.5 models. Case (a) stands for the zooming evaluation, where models successfully identify the view containing useful information and generate the correct result. Case (b) illustrates a failure in the Shift-R evaluation, where all the models continue shifting to new views until all views are used. Though including the correct views, the additional views severely distract the reasoning process, where three out of four employed models produce incorrect answers. To explore how human-like mixed evaluation affects the visual reasoning process, we further exam this failure case using GPT-4o. As shown in Figure 5 case (c), GPT-4o first zooms into the “upper left” and “upper right” views, then discards the “upper right” view and shifts to the “lower left” one, which finally leads to the correct answer. Notably, in the final preserved views, distracting information (the highest price tag on papaya, “69”) is screened out. This indicates that GPT-4o exhibits decent active perception abilities to move the field of view, locate details, and filter out distracting information.
I.2 Cases of Giving Human-annotated Clues